# Giulia Gaggero and Elisa Gorla

### v2

@@ 3 Random polynomial systems
Let $\mathbb{F}$ be a field and let $\overline{\mathbb{F}}$ be its algebraic closure. Denote by $\mathbb{P}^{t}$ the $t$-dimensional projective space over $\overline{\mathbb{F}}$. Let $d_{1},\ldots,d_{m}$ be positive integers and let $\mathcal{U}\subseteq\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$ be a nonempty Zariski-open set. For a homogeneous polynomial $f$, denote by $[f]\in\mathbb{P}^{\binom{n+\deg(f)-1}{n-1}-1}$ the projective point, whose coordinates are the coefficients of $f$. A homogeneous system =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ with $\deg(f_{i})=d_{i}$ for $1\leq i\leq m$ is generic with respect to $\mathcal{U}$ if $([f_{1}],\ldots,[f_{m}])\in\mathcal{U}$. It is generic if it is generic with respect to $\mathcal{U}$, for some nonempty Zariski-open set $\mathcal{U}$. An arbitrary system $\mathcal{F}$ is generic or generic with respect to $\mathcal{U}$ if $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is.
In the cryptographic literature, random sequences of polynomials are often assumed to be cryptographic semiregular sequences, see e.g. [2, 4]. This is the case in the cryptoanalysis of several systems, as e.g. [15]. The next definition appears in [2, Definition 3.2.1 and Definition 3.2.4][5, Definition 5], and [7, Definition 5 and Definition 9].
@@
Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a homogeneous system.
If $\mathbb{F}\neq\mathbb{F}_{2}$, we say that $f_{1},\ldots,f_{m}$ are a cryptographic semiregular sequence if for all $1\leq i\leq m$ and all $g_{i}\in R$ such that $g_{i}f_{i}\in(f_{1},\ldots,f_{i-1})$ and $\deg(g_{i}f_{i})<d_{\operatorname{reg}}(\mathcal{F})$, one has that $g_{i}\in(f_{1},\ldots,f_{i-1})$.
If $\mathbb{F}=\mathbb{F}_{2}$, we say that $f_{1},\ldots,f_{m}\in R/(x_{1}^{2},\ldots,x_{n}^{2})$ are a cryptographic semiregular sequence if for all $1\leq i\leq m$ and all $g_{i}\in R/(x_{1}^{2},\ldots,x_{n}^{2})$ such that $g_{i}f_{i}\in(f_{1},\ldots,f_{i-1})$ and $\deg(g_{i}f_{i})<d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{2},\ldots,x_{n}^{2}\})$, one has that $g_{i}\in(f_{1},\ldots,f_{i})$.
Arbitrary polynomials $f_{1},\ldots,f_{m}$ are a cryptographic semiregular sequence if $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are a cryptographic semiregular sequence.
In this paper, we use the word cryptographic semiregular sequence in order to distinguish the concept of semiregularity used in the cryptographic literature from the concept of semiregularity originally introduced by Pardue [24, 25], which inspired it. The original definition by Pardue is given over an infinite field $\mathbb{F}$. As we are interested also in dealing with finite fields, we extend it in the natural way.
@@
Let $\mathbb{F}$ be an infinite field and let $R=\mathbb{F}[x_{1},\ldots,x_{n}]$. Let $I$ be a homogeneous ideal and let $A=R/I$. A polynomial $f\in R_{d}$ is semiregular on $A$ if for every $e\geq d$, the vector space map $A_{e-d}\rightarrow A_{e}$ given by multiplication by $f$ has maximal rank (that is, it is either injective or surjective). If $\mathbb{F}$ is a finite field, let $\overline{R}=\overline{\mathbb{F}}[x_{1},\ldots,x_{n}]$. Then $f$ is semiregular on $A$ if it is semiregular on $\overline{R}/I\overline{R}$.
A sequence of homogeneous polynomials $f_{1},\ldots,f_{m}$ is a semiregular sequence if $f_{i}$ is semiregular on $A/(f_{1},\ldots,f_{i-1})$ for all $1\leq i\leq m$.
It follows from [25, Proposition 1] that, if $\mathbb{F}\neq\mathbb{F}_{2}$, then a semiregular sequence is also a cryptographic semiregular sequence. The converse does not hold, as shown in [25], see the example just below [25, Proposition 1].
@@
Let $q>2$ and let $f_{1}=x_{1}^{q}-x_{1},\ldots,f_{n}=x_{n}^{q}-x_{n},f_{n+1}=f\in R=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$. The sequence $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular if and only if the sequence $f_{1}^{\operatorname{\mathrm{top}}}=x_{1}^{q},\ldots,$ $f_{n}^{\operatorname{\mathrm{top}}}=x_{n}^{q},f_{n+1}^{\operatorname{\mathrm{top}}}=f^{\operatorname{\mathrm{top}}}$ is semiregular.
@@
It suffices to show that, if the sequence $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular, then the sequence $x_{1}^{q},\ldots,x_{n}^{q},f^{\operatorname{\mathrm{top}}}$ is semiregular. We start by observing that the sequence $x_{1}^{q},\ldots,x_{n}^{q}$ is regular. Therefore, in order to show that $x_{1}^{q},\ldots,x_{n}^{q},f^{\operatorname{\mathrm{top}}}$ is semiregular, it suffices to show that $f^{\operatorname{\mathrm{top}}}$ is semiregular on $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{q},\ldots,x_{n}^{q})$. If $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular and $d=\deg(f)$, then the Hilbert series of $R/(\mathcal{F}^{\operatorname{\mathrm{top}}})$ is $[(1-z^{d})(1+z+\ldots+z^{q-1})^{n}]$ by [7, Proposition 6]. Here for $p(z)=\sum_{i=0}^{\infty}p_{i}z^{i}\in\mathbb{Z}[[z]]$, we denote by $\delta(p)=\min\{i\geq 0\mid p_{i}\leq 0\}-1$ and define $[p(z)]=\sum_{i=0}^{\delta(p)}p_{i}z^{i}$. Then $f^{\operatorname{\mathrm{top}}}$ is semiregular on $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{q},\ldots,x_{n}^{q})$ and $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a semiregular sequence by [25, Proposition 1]. ∎
Pardue in [24, 25] shows that Fröberg’s Conjecture [22], a conjecture which has attracted a lot of attention within the commutative algebra community and that is widely believed to hold, is equivalent to the following
@@
In other words, Fröberg conjectures that the set of semiregular sequences of polynomials of given degrees contains a dense Zarisky-open set. If the conjecture is true, then a sequence of polynomials of given degrees is semiregular with high probability, provided that the ground field has large enough cardinality. It follows that, if $\mathbb{F}=\mathbb{F}_{q}$ with $q\gg 0$ and Fröberg’s Conjecture holds, then a sequence of polynomials of given degrees is a cryptographic semiregular sequence with high probability. In addition, most cryptographic semiregular sequences are also semiregular sequences, as the set of semiregular sequences conjecturally contains a dense open set.
In [2, Section 3.2], Bardet conjectures that a sequence of polynomials with coefficients in $\mathbb{F}_{2}$ is cryptographic semiregular with high probability. This conjecture is motivated by experimental evidence, see also [4, Conjecture 2]. The conjecture was later disproved by Hodges, Molina, and Schlather, who in [23] prove that there are choices of the parameters for which no cryptographic semiregular sequence exists over $\mathbb{F}_{2}$. This is the case, e.g., for $m=1$ and $n>3d_{1}$. In the sequel, we propose a notion of randomness which applies to any choice of the system parameters.
In this paper, we propose two dense Zariski-open sets, which can be used to formalize the intuitive idea of a random system. The set $\mathcal{V}$ corresponds to systems of $m$ polynomials in $n$ variables which contain a regular sequence of length $n$, while the set $\mathcal{U}$ parametrizes systems of $m$ polynomials which contain a regular sequence of $n$ polynomials of the smallest possible degrees. Notice that $\mathcal{V}$ contains $\mathcal{U}$ and, if $m\geq n$, it also contains the set of cryptographic semiregular sequences and that of semiregular sequences.
@@
Fix $m\geq n\geq 1$ and $1\leq d_{1}\leq\ldots\leq d_{m}$. For any multiset $\Delta$ of cardinality $n$ contained in the multiset $\{d_{1},\ldots,d_{m}\}$, let $\mathcal{U}_{\Delta}$ be the subset of $\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$ whose points correspond to polynomials $f_{1},\ldots,f_{m}\in R$ such that $(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}})$ contains a regular sequence in the degrees of $\Delta$. Let
$\mathcal{U}=\mathcal{U}_{\{d_{1},\ldots,d_{n}\}}\;\;\;\mbox{ and }\;\;\;\mathcal{V}=\bigcup_{\Delta\subseteq\{d_{1},\ldots,d_{m}\},|\Delta|=n}\mathcal{U}_{\Delta}.$
An algebraically random system of $m$ polynomials of degrees $d_{1},\ldots,d_{m}$ in $n$ variables is a system =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ such that $\deg(f_{i})=d_{i}$ for all $1\leq i\leq m$ and $([f_{1}],\ldots,[f_{m}])\in\mathcal{U}$.
It is well-known that $\mathcal{U}_{\{d_{1},\ldots,d_{n}\}}$ is a dense open set for any choice of $1\leq d_{1}\leq\ldots\leq d_{n}$, if $m=n$. This implies that any $\mathcal{U}_{\Delta}$ as in Definition 12 is a dense open set, hence also $\mathcal{U}$ and $\mathcal{V}$ are. Notice that the degree of regularity is finite for all systems in $\mathcal{U}_{\Delta}$ and for any choice of $\Delta$, therefore it is also finite for the systems in $\mathcal{U}$ and $\mathcal{V}$. In particular, the degree of regularity of an algebraically random system is finite. Notice moreover that, if $\mathcal{F}$ is homogeneous, then $\mathcal{V}$ is the set of polynomial systems of degrees $d_{1},\ldots,d_{m}$ for which the degree of regularity is finite.
@@
Notice moreover that, if $m=n$, an algebraically random system is a system for which $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a regular sequence. Since this case is well-studied in the cryptographic literature, in the sequel we often assume $m>n$.

### v3

@@ 3 Systems that contain a regular sequence in a given degree
Let $\mathbb{F}$ be a field, and let $\overline{\mathbb{F}}$ be its algebraic closure. Denote by $\mathbb{P}^{t}$ the $t$-dimensional projective space over $\overline{\mathbb{F}}$. Let $d_{1},\ldots,d_{m}$ be positive integers. For a homogeneous polynomial $f$, denote by $f]\in\mathbb{P}^{\binom{n+\deg(f)-1}{n-1}-1}$ the projective point whose coordinates are the coefficients of $f$. A homogeneous system $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ with $\deg(f_{i})=d_{i}$ for $1\leq i\leq m$ corresponds to a point $([f_{1}],\ldots,[f_{m}])\in\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$.
A random system of degrees $d_{1},\ldots,d_{m}$ over a field $\mathbb{F}$ is a system of polynomials of the given degrees whose coefficients are chosen uniformly at random in the given field. Similarly, a random homogeneous system of degrees $d_{1},\ldots,d_{m}$ is a system of homogeneous polynomials of the given degrees, whose coefficients are chosen uniformly at random. The concept of random system is used extensively in the cryptographic literature, where random instances of structured polynomial systems are often assumed to behave as random systems.
Over an infinite field, a generic property in the sense of algebraic geometry is by definition a property which holds over a nonempty Zarisky-open set. We refer to [16, Section 1.3] for a short reminder on the Zarisky topology. Since every nonempty Zarisky-open set is dense, the property holds for “almost every system”. Therefore, over an infinite field, genericity is the algebraic-geometric analog of being true with high probability. Over a finite field $\mathbb{F}_{q}$, however, estimating the probability that a system has a given property amounts to estimating the number of $\mathbb{F}_{q}$-rational points of the open set that corresponds to the property over the algebraic closure, a task which is often far from trivial. Nevertheless, if $q$ is large enough (or if we consider a finite extension of $\mathbb{F}_{q}$ of large enough cardinality), a random homogeneous system belongs to any given nonempty Zarisky-open set with high probability. For the convenience of the reader, we show how to deduce this from known results. For simplicity, we only state the argument for one projective space, but the same argument can be applied to a product of projective spaces.
Proposition 6.
Let $\mathbb{P}^{t}$ be the $t$-dimensional projective space over $\bar{\mathbb{F}_{q}}$. Let $\mathcal{U}\subseteq\mathbb{P}^{t}$ be a nonempty Zarisky-open set defined over $\mathbb{F}_{q}$. If $q$ is sufficiently large, then a randomly chosen $\mathbb{F}_{q}$-rational point $P\in\mathbb{P}^{t}$ belongs to $\mathcal{U}$ with high probability.
Proof.
Let $\mathcal{X}$ be the set-theoretic complement of $\mathcal{U}$ in $\mathbb{P}^{t}$. Let $Q$ be a power of $q$. It suffices to show that the probability that a randomly chosen $\mathbb{F}_{Q}$-rational point belongs to $\mathcal{X}$ tends to zero as $Q$ tends to infinity. Since $\mathcal{X}$ is a projective algebraic variety, it follows from [25, Theorem 3.1] that the number $|\mathcal{X}(\mathbb{F}_{Q})|$ of $\mathbb{F}_{Q}$-rational points of $\mathcal{X}$ is bounded from above by a polynomial $p(Q)$, whose coefficients depend on the degrees and dimensions of the irreducible components of $\mathcal{X}$. In particular,
$\deg p(Q)\leq\max\{\dim(\mathcal{X}),2\dim(\mathcal{X})-t\}\leq t-1.$
Since the number of $\mathbb{F}_{Q}$-rational points of $\mathbb{P}^{t}$ is $\sum_{i=0}^{t}Q^{i}$, then the probability that an $\mathbb{F}_{Q}$-rational belongs to $\mathcal{X}$ is bounded from above by
[…]
Hence, the probability tends to zero as $Q$ tends to infinity. ∎
In the cryptographic literature, random polynomial systems are often assumed to be cryptographic semiregular sequences, see e.g. [3, 6]. This is the case in the cryptoanalysis of several systems, as e.g. [19]. The next definition appears in [3, Definition 3.2.1 and Definition 3.2.4], [7, Definition 5], and [9, Definition 5 and Definition 9].
@@
Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a homogeneous system.
If $\mathbb{F}\neq\mathbb{F}_{2}$, we say that $f_{1},\ldots,f_{m}$ are a cryptographic semiregular sequence if for all $1\leq i\leq m$ and all $g_{i}\in R$ such that $g_{i}f_{i}\in(f_{1},\ldots,f_{i-1})$ and $\deg(g_{i}f_{i})<d_{\operatorname{reg}}(\mathcal{F})$, we have $g_{i}\in(f_{1},\ldots,f_{i-1})$.
If $\mathbb{F}=\mathbb{F}_{2}$, we say that $f_{1},\ldots,f_{m}\in R/(x_{1}^{2},\ldots,x_{n}^{2})$ are a cryptographic semiregular sequence if for all $1\leq i\leq m$ and all $g_{i}\in R/(x_{1}^{2},\ldots,x_{n}^{2})$ such that $g_{i}f_{i}\in(f_{1},\ldots,f_{i-1})$ and $\deg(g_{i}f_{i})<d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{2},\ldots,x_{n}^{2}\})$, we have $g_{i}\in(f_{1},\ldots,f_{i})$.
A sequence of polynomials (not necessarily homogeneous) $f_{1},\ldots,f_{m}$ is a cryptographic semiregular sequence if $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are a cryptographic semiregular sequence.
In this paper, we use the expression cryptographic semiregular sequence in order to distinguish the concept of semiregularity used in the cryptographic literature from the concept of semiregularity originally introduced by Pardue [35, 36], which inspired it. The original definition by Pardue is given over an infinite field $\mathbb{F}$. As we are interested in dealing with finite fields, we extend it to an arbitrary field in the natural way.
@@
Let $\mathbb{F}$ be a field, and let $R=\mathbb{F}[x_{1},\ldots,x_{n}]$. Let $I$ be a homogeneous ideal, and let $A=R/I$. A polynomial $f\in R_{d}$ is semiregular on $A$ if for every $e\geq d$, the vector space map $A_{e-d}\rightarrow A_{e}$ given by multiplication by $f$ has maximal rank (that is, it is either injective or surjective). A sequence of homogeneous polynomials $f_{1},\ldots,f_{m}$ is a semiregular sequence if $f_{i}$ is semiregular on $A/(f_{1},\ldots,f_{i-1})$ for all $1\leq i\leq m$.
It follows from [36, Proposition 1] that, if $\mathbb{F}\neq\mathbb{F}_{2}$, then a semiregular sequence is also a cryptographic semiregular sequence. The converse does not hold, as shown in [36], see the example just below [36, Proposition 1].
@@
Let $q>2$ and let $f_{1},\ldots,f_{n+1}\in R=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ be polynomials of degree $d_{1},\ldots,d_{n+1}$, respectively. Assume that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n}^{\operatorname{\mathrm{top}}}$ is a regular sequence. The sequence $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular if and only if the sequence $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n+1}^{\operatorname{\mathrm{top}}}$ is semiregular. In particular, the sequence $x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n},f$ is cryptographic semiregular if and only if the sequence $x_{1}^{q},\ldots,x_{n}^{q},f^{\operatorname{\mathrm{top}}}$ is semiregular.
@@
Let $\mathcal{F}=\{f_{1},\ldots,f_{n+1}\}$. It suffices to show that if the sequence $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular, then the sequence $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n+1}^{\operatorname{\mathrm{top}}}$ is semiregular. Since $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n}^{\operatorname{\mathrm{top}}}$ is a regular sequence, it suffices to show that $f_{n+1}^{\operatorname{\mathrm{top}}}$ is semiregular on $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n}^{\operatorname{\mathrm{top}}})$. If $f_{1},\ldots,f_{n+1}$ is cryptographic semiregular, then by [9, Proposition 6] the Hilbert series of $R/(\mathcal{F}^{\operatorname{\mathrm{top}}})$ is $(1-z^{d_{n+1}})\prod_{i=1}^{n}(1+z+\ldots+z^{d_{i}-1})]$. Here, for $p(z)=\sum_{i=0}^{\infty}p_{i}z^{i}\in\mathbb{Z}[[z]]$, we denote by $\delta(p)=\min\{i\geq 0\mid p_{i}\leq 0\}-1$ and define $p(z)]=\sum_{i=0}^{\delta(p)}p_{i}z^{i}$. Then $f_{n+1}^{\operatorname{\mathrm{top}}}$ is semiregular on $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n}^{\operatorname{\mathrm{top}}})$ and $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a semiregular sequence by [36, Proposition 1]. ∎
Pardue in [35, 36] shows that Fröberg’s Conjecture [30], a conjecture which has attracted a lot of attention within the commutative algebra community and is widely believed to hold, is equivalent to the following
@@
In other words, Fröberg’s Conjecture is equivalent to the statement that the set of semiregular sequences of polynomials of given degrees contains a dense Zarisky-open set. If the conjecture is true, then a sequence of polynomials of given degrees is semiregular with high probability, provided that the ground field has large enough cardinality. It follows that, if $\mathbb{F}=\mathbb{F}_{q}$ with $q\gg 0$ and Fröberg’s Conjecture holds, then a sequence of polynomials of given degrees is a cryptographic semiregular sequence with high probability. In addition, for $q\gg 0$ most cryptographic semiregular sequences are also semiregular sequences, as the set of semiregular sequences conjecturally contains a dense open set.
In [3, Section 3.2], Bardet conjectures that a sequence of polynomials with coefficients in $\mathbb{F}_{2}$ is cryptographic semiregular with high probability. This conjecture is motivated by experimental evidence, see also [6, Conjecture 2]. The conjecture was later disproved by Hodges, Molina, and Schlather, who in [31] prove that there are choices of the parameters for which no cryptographic semiregular sequence exists over $\mathbb{F}_{2}$. This is the case, e.g., for $m=1$ and $n>3d_{1}$.
In this paper, we study the Zariski-open set consisting of systems of $m$ polynomials in $n$ variables which generate an ideal that contains a regular sequence of $n$ polynomials in a given degree $D$.
@@
Let $\mathbb{F}$ be a field and let $\overline{\mathbb{F}}$ be its algebraic closure. Denote by $\mathbb{P}^{t}$ the $t$-dimensional projective space over $\overline{\mathbb{F}}$. Fix $m\geq n\geq 1$ and $1\leq d_{1}\leq\ldots\leq d_{m}$. Let $\mathcal{U}_{D}\subseteq\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$ be the set of $([g_{1}],\ldots,[g_{m}])\in\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$ such that the ideal $(g_{1},\ldots,g_{m})\subseteq\overline{\mathbb{F}}[x_{1},\ldots,x_{n}]$ contains a regular sequence of $n$ polynomials of degree $D$. We say that the system $\mathcal{F}$ is regular in degree $D$ if $([f_{1}^{\operatorname{\mathrm{top}}}],\ldots,[f_{m}^{\operatorname{\mathrm{top}}}])\in\mathcal{U}_{D}$.
Fix $m\geq n\geq 1$, $1\leq d_{1}\leq\ldots\leq d_{m}$, and $\mathbb{F}$ a field. If $D<d_{n}$, then $\mathcal{U}_{D}=\emptyset$. In fact, let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ with $\deg(f_{i})=d_{i}$. Then $(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}})_{D}=(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{n-1}^{\operatorname{\mathrm{top}}})_{D}$ cannot contain a regular sequence of $n$ polynomials. Therefore, the system $\mathcal{F}$ can be regular in degree $D$, only if $D\geq d_{n}$.
The next proposition provides some useful facts connected to the property of being regular in a given degree. The statement follows from standard facts from commutative algebra. Recall that a system defined over a field $\mathbb{F}$ is zero-dimensional if it has finitely many solutions over $\overline{\mathbb{F}}$.
Proposition 12.
Let $\mathbb{F}$ be a field, and let $m\geq n\geq 1$ and $1\leq d_{1}\leq\ldots\leq d_{m}$. Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ with $\deg(f_{i})=d_{i}$. The following are equivalent:
  1. 1.
    $d_{\operatorname{reg}}(\mathcal{F})<+\infty$,
  2. 2.
    $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is zero-dimensional,
  3. 3.
    $\mathcal{F}^{\operatorname{\mathrm{top}}}\in\mathcal{U}_{D}$ for some $D\geq 0$,
  4. 4.
    $\mathcal{F}^{\operatorname{\mathrm{top}}}\in\mathcal{U}_{D}$ for all $D\geq d_{m}$.
In particular, if $d_{1}=\ldots=d_{m}=D$ and $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is zero-dimensional, then $\mathcal{F}$ is regular in degree $D$.
@@
It follows from Proposition 12 that, if $m=n$, a system $\mathcal{F}$ has finite degree of regularity if and only if $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a regular sequence. Since this case is well-studied, in the sequel we often focus on the case where $m>n$.
Notice moreover that, if $\mathcal{F}$ is cryptographic semiregular, then it satisfies the equivalent conditions of Proposition 12. In other words, every cryptographic semiregular system is regular in degree $D$ for every $D\geq d_{m}$ (and possibly in smaller degree). However, the converse does not hold.
The statement that every cryptographic semiregular sequence is regular in some degree also implies that being regular in degree $D$ is a generic property, conditionally to Conjecture 10. More precisely, Proposition 12 implies that, if $D\geq d_{n}$, then $\mathcal{U}_{D}$ contains a dense open set of $\mathbb{P}^{\binom{n+d_{1}-1}{n-1}-1}\times\ldots\times\mathbb{P}^{\binom{n+d_{m}-1}{n-1}-1}$.
Even over a finite field, one advantage of considering systems which are regular in degree $D$ is that, unlike cryptographic semiregular systems [31], they exist over every field and for any choice of $m\geq n$ and $D\geq d_{n}$. In other words, if the set $\mathcal{U}_{D}$ is nonempty over $\overline{\mathbb{F}_{q}}$, then it always contains $\mathbb{F}_{q}$-rational points.
Example 13.
The set $\mathcal{U}_{D}$ contains $\mathbb{F}_{q}$-rational points for any choice of $n,m,q$ and $1\leq d_{1}\leq\ldots\leq d_{m}$ with $n\leq m$ and $D\geq d_{n}$. This follows from the existence of homogeneous regular sequences of any given degrees in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$. Some homogeneous regular sequences in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of degrees $d_{1},\ldots,d_{n}$ are for example
$x_{1}^{d_{1}}+g_{1},x_{2}^{d_{2}}+g_{2},\ldots,x_{n}^{d_{n}}+g_{n},$
where $g_{i}\in\mathbb{F}_{q}[x_{i+1},\ldots,x_{n}]_{d_{i}}$. It follows that the system
$\{x_{1}^{d_{1}}+g_{1},x_{2}^{d_{2}}+g_{2},\ldots,x_{n}^{d_{n}}+g_{n},f_{n+1},\ldots,f_{m}\}$
is regular in degree $D$ for every $D\geq d_{n}$ and for any $f_{n+1},\ldots,f_{m}$ of the appropriate degrees.
We stress that Proposition 12 applies to most zero-dimensional systems, since for most zero-dimensional systems over an algebraically closed field, the top degree part of the system $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is also zero-dimensional. This statement can be formalized using the algebraic geometric language of dense open sets. Nevertheless, there exist zero-dimensional systems $\mathcal{F}$ such that $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is not zero-dimensional.
Example 14.
Let $f_{1}=x^{3}$, $f_{2}=xy^{2}+y^{2}$, $f_{3}=x^{2}y$. The system is zero-dimensional, as $\sqrt{(f_{1},f_{2},f_{3})}=(x,y)$. However, the system is not regular in degree $3$, since $(f_{1}^{\operatorname{\mathrm{top}}},f_{2}^{\operatorname{\mathrm{top}}},f_{3}^{\operatorname{\mathrm{top}}})=(x^{3},x^{2}y,xy^{2})$ does not contain a regular sequence in degree $3$. In fact, $(f_{1}^{\operatorname{\mathrm{top}}},f_{2}^{\operatorname{\mathrm{top}}},f_{3}^{\operatorname{\mathrm{top}}})$ is not even zero-dimensional, as $\sqrt{(f_{1}^{\operatorname{\mathrm{top}}},f_{2}^{\operatorname{\mathrm{top}}},f_{3}^{\operatorname{\mathrm{top}}})}=(x)$.
Example 15.
The system $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}$ is regular in degree $q$, for any system $\mathcal{F}$. In fact, $(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}},x_{1}^{q},\ldots,x_{n}^{q})_{q}$ contains the regular sequence $x_{1}^{q},\ldots,x_{n}^{q}$.
4 Bounds on the degree of regularity
Throughout the section, $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ denotes a system of $m$ polynomials in $n$ variables. We always assume that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent and $m\geq n$. Notice that
$d_{\operatorname{reg}}(\mathcal{F})\geq d_{n}$
by definition. In this section, we establish an upper bound for the degree of regularity of a system $\mathcal{F}$ of polynomials of equal degree $D$, under the assumption that $\mathcal{F}$ is regular in degree $D$. We also prove an upper bound for the degree of regularity of a system $\mathcal{F}$ of polynomials of equal degree $D$, for $D\geq q$. In combination with Theorem 3 and Theorem 4, these results provide us with upper bounds for the solving degree of $\mathcal{F}$.